What line goes through (-9, 0) that has a slope of 3/5? Write your answer in standard form Ax + By = C.

Answer by checkley71(8403)

0=3/5*-9+b
0=-29/5+b
b=29/5 THUS THE LINE EQUATION IS
Y=3X/5+29/5 OR -3X/5+Y=29/5